1.Solve by substitution or elimination method: Thanks!
4x – 5y = 14
-12x + 15y = -42

Answer by Alan3354(69443)

Solve by substitution or elimination method: Thanks!
4x – 5y = 14
-12x + 15y = -42
------------
The 2nd eqn is the 1st multiplied by -3 --> dependent.
No solution.